# DeepEdit: Knowledge Editing as Decoding with Constraints

### v2

@@ 4 New Benchmarks for More Precise and Challenging Evaluation
We provide two new benchmarks for the evaluation of KE methods: MQuAKE-2002 and MQuAKE-hard. Both are built from the MQuAKE Zhong et al. (2023), a recent KE benchmark that is designed to evaluate the KE methods on helping LLMs to answer multi-hop questions given new knowledge. Every instance in MQuAKE includes a multi-hop question and one to four related edited facts, every of which can alter the ground-truth answer to the question. Zhong et al. (2023) suggests using a randomly sampled 3,000 instance subset of MQuAKE to do the evaluation, which reduces the experimental costs. This evaluation benchmark is termed as MQuAKE-3k.
4.1 Annotation Mistakes in One-Third Instances of MQuAKE-3k
@@
There are two issues of using MQuAKE-3k to evaluate KE. The first issue is that the new knowledge from different instances MQuAKE-3k can cause conflicts and mistakes to the ground-truth answers. In other words, the ground-truth answer from instance A can be altered by the new knowledge from another instance B. We show an example of knowledge conflicts in MQuAKE-3k in Figure 7. These knowledge conflicts will make the ground-truth answers incorrect given the new knowledge that are conflicted with the answers because the inference on every instance would retrieve the new knowledge from all instances. We observe that 998 instances’ ground-truth answers are broken by the new knowledge from other instances.
@@ 4.2 New Benchmark MQuAKE-2002 for More Precise Evaluation
To address the issue of annotation mistakes in MQuAKE-3k, we provide a new subset of MQuAKE-3k, which does not have any knowledge conflict across instances. This subset includes 2,002 instances, so we term it as MQuAKE-2002. We filter out the instances of which the ground-truth answers are broken by the new knowledge from other instances to produce MQuAKE-2002. Compared with MQuAKE-3k, our MQuAKE-2002 provides a more precise evaluation for knowledge editing methods, since it remove the annotation mistakes due to knowledge conflicts across instances. The data statistics of MQuAKE-2002 is provided in Table 1.
4.3 New Benchmark MQuAKE-hard for More Challenging Evaluation
The second issue of MQuAKE-3k is that more than 66% instances in MQuAKE-3k only contain at most two edited facts that influence the answers, which are not challenging enough to evaluate the knowledge editing methods on handling multiple edited facts that can alter the ground-truth reasons. we construct a more challenging subset of MQuAKE by selecting the instances that contain the highest number of edited facts per instance. We term this challenging set as MQuAKE-hard, which includes 429 instances and every instance contains four edited facts. MQuAKE-hard has no overlap with MQuAKE-3k. We term this challenging set as MQuAKE-hard, which includes 429 instances and every instance contains four edited facts. MQuAKE-hard has no overlap with MQuAKE-3k. The data statistics of MQuAKE-hard is also provided in Table 1.
Benchmark # Instances # Hops per Instance # Edited Facts per Instance # Conflicted Instances MQuAKE-3k Zhong et al. (2023) 3,000 3.0 2.0 998 MQuAKE-2002 2,002 2.7 2.2 0 MQuAKE-hard 4,29 4.0 4.0 0
@@ 5 Experiments
In this section, we evaluate the KE performance of our DeepEdit method when applied to LLMs, and compare against a strong baseline MeLLo Zhong et al. (2023). Our experimental settings closely follow those of the previous work Zhong et al. (2023) to ensure a fair comparison.
@@ 5.1 Experimental Settings
We take the following KE benchmarks MQuAKE-3k Zhong et al. (2023), MQuAKE-2002, and MQuAKE-hard for evaluation. The introduction of the above benchmarks are provided in Section 4. We follow the existing work Zhong et al. (2023) to use accuracy as the main evaluation metric. For every instance, we test the accuracy of the first question in the instance, since different questions in the same instance are semantically equivalent and correspond to the same ground-truth answer. We follow Zhong et al. (2023) to set the KE batch size as 100 and full batch for evaluation. The KE batch size means the number of instances that provide new knowledge for retrieval.
@@
Compared methods. We take the strong KE methods into comparison: MEND Mitchell et al. (2021), MEMIT Meng et al. (2022b), IKE Zheng et al. (2023a), ICE Cohen et al. (2023), and MeLLo Zhong et al. (2023). MEND produces weight updates by transforming the raw fine-tuning gradients given an edited fact. MEMIT updates feedforward networks in a range of layers to encode the edited facts. IKE and ICE conduct KE by adding new knowledge to the input prompts. MeLLo is the state-of-the-art KE method that prompts the LLMs to generate subquestions and append new knowledge to the prompts. The former two KE methods are only applicable to white-box LLMs, while the latter three can fit the block-box LLMs.
@@
Method MQuAKE-3k MQuAKE-hard batch size of 100 full batch batch size of 100 full batch llama2-7b w/ MEND Mitchell et al. (2021) 7.2 3.9 0.5 0.2 llama2-7b w/ MEMIT Meng et al. (2022b) 7.9 4.2 0.9 0.5 llama2-7b w/ IKE Zheng et al. (2023a) 12.5 6.2 1.2 0.5 llama2-7b w/ ICE Cohen et al. (2023) 12.9 6.3 1.4 0.7 llama2-7b w/ MeLLo Zhong et al. (2023) 16.3 10.8 1.9 1.6 llama2-7b w/ DeepEdit (Ours) 27.6 11.2 14.5 7.0 gpt-3.5 w/ IKE Zheng et al. (2023a) 15.5 8.3 1.6 1.2 gpt-3.5 w/ ICE Cohen et al. (2023) 16.1 8.5 2.1 1.4 gpt-3.5 w/ MeLLo Zhong et al. (2023) 29.9 20.0 3.7 1.6 gpt-3.5 w/ DeepEdit (Ours) 47.2 38.0 51.3 48.0 text-davinci-003 w/ IKE Zheng et al. (2023a) 18.2 8.9 1.9 1.2 text-davinci-003 w/ ICE Cohen et al. (2023) 19.0 8.6 2.3 1.2 text-davinci-003 w/ MeLLo Zhong et al. (2023) 32.2 25.6 4.7 2.3 text-davinci-003 w/ DeepEdit (Ours) 49.1 38.8 52.0 49.2
@@
Table 2: Experimental results (accuracy; %) on the dataset MQuAKE-3k and MQuAKE-hard. The batch size means the number of instances providing the edited facts for knowledge retrieval. gpt-3.5 stands for the model gpt-3.5-turbo-instruct. The best KE result on every LLM is highlighted in bold font.
Method MQuAKE-2002 gpt-3.5 w/ MeLLo 25.1 gpt-3.5 w/ DeepEdit 53.7
Table 3: Experimental results (accuracy; %) of the full-batch knowledge editing on the benchmark MQuAKE-2002.
Verifier [Conciseness] [Coherence] [Receptiveness] llama2-7b 82.3 77.9 86.3 gpt-3.5 92.8 91.5 92.2
@@
Table 4: Experimental results (verification accuracy; %) of our verifiers for different constraints on the dataset MQuAKE-3k.
gpt-3.5 w/ DeepEdit 49.1 $\Delta$ w/o [Conciseness] 46.1 $\downarrow 6\%$ w/o [Coherence] 42.9 $\downarrow 13\%$ w/o [Receptiveness] 45.2 $\downarrow 8\%$ w/o [Pertinence] 46.6 $\downarrow 5\%$
Table 5: Experimental results (accuracy; %) on the dataset MQuAKE-3k of DeepEdit with KE batch size of 100.
@@
Model configuration. We set the hyper-parameters of the baseline methods as suggested by their papers. We apply the KE methods to the strong LLM text-davinci-003, gpt-3.5-turbo-instruct (short as gpt-3.5), and llama2-7b (short as llama2) OpenAI (2022); Touvron et al. (2023) for evaluation. The former two are black-box LLMs and the latter one is a white-box LLM. By default, we set the decoding temperature as 0.0 to minimize the randomness of LLMs’ outputs. The DeepEdit’s verifiers use the same LLM as the generator.
@@
5.2 Overall Performance
We incorporate DeepEdit with several choices of LLMs to test its effectiveness on editing LLMs’ knowledge. We report the test accuracy on MQuAKE-3k and MQuAKE-hard in Table 2. When the KE batch size is 100, our DeepEdit improves the accuracy of question answering with KE by 69%, 58%, and 52% over MeLLo on the LLMs llama2, gpt-3.5-turbo-instruct, and text-davinci-003 respectively over the strong baseline method MeLLo on MQuAKE-3k. On MQuAKE-hard, our DeepEdit outperforms the strong baseline method MeLLo by $\times$7, $\times$12, and $\times$10 times in terms of question answering accuracy with KE on the LLMs llama2, gpt-3.5, and text-davinci-003 respectively.
The experimental results on MQuAKE-2002 is shown in Table 3. When the KE batch size is full batch, our DeepEdit still outperforms the baseline methods and improves the KE accuracy by a significant margin. Our DeepEdit enhances the LLMs’ reasoning on new knowledge through higher conciseness, coherence, pertinence, and receptiveness. The above advantages are especially important when the LLMs have to utilize multiple new knowledge to get the correct answers, since more new knowledge requires the LLMs to be correctly aware of the new knowledge on more reasoning steps. The correct knowledge incorporation with our decoding constraints significantly improves the KE performance given more new knowledge. Overall, our DeepEdit achieves substantial improvements on KE for LLMs and consistently outperforms the baseline methods on different benchmarks.
5.3 Analysis on Verifiers
We conduct deep analysis on the different verifiers of DeepEdit. We first test the accuracy of different verifiers to verify the corresponding constraint. We randomly sample 4,389 pairs of reasoning steps from the ground-truth reasoning steps of MQuAKE as the positive pairs that satisfy the constraints and randomly pairing the same number of steps from different questions as the negative examples. We evaluate different verifiers based on llama2 and gpt-3.5 on the sampled reasoning step pairs. We report the verification accuracy in Table 4. We observe high verification accuracy of around 80% and over 90% on different constraints achieved by llama-2 and gpt-3.5 respectively. Also, we observe that the performance of verifiers based on llama2 and gpt-3.5 are positively related to the LLMs’ performance with our DeepEdit, since the effective constraint assessment forms the basis of effective reasoning of LLMs with our DeepEdit.
5.4 Ablation Study
We conduct ablation studies on DeepEdit to empirically examine the contribution of its main technical components. We evaluate the influence of different constraints on DeepEdit. We report the experimental results of the ablation study of constraints in Table 5. We observe that removing each constraint leads to substantial performance degradation in different degrees. This indicates that different constraints contribute to more effective KE from the perspectives of progress, relevance, coherence, and awareness as a whole, and removing any of them will lead to more mistakes in the reasoning of LLMs. The above results also validate that the LLMs’ outputs cannot automatically satisfy the constraints that represent the desired properties without our DeepEdit.
5.5 Case Study

### v3

@@ 4 New Benchmarks for More Precise and Challenging Evaluation of KE
We provide two new benchmarks for the evaluation of KE methods: MQuAKE-2002 and MQuAKE-hard. Both are built based on MQuAKE Zhong et al. (2023), a recent KE dataset that is designed to evaluate the KE methods on helping LLMs to answer multi-hop questions given new knowledge. Every instance in MQuAKE includes a multi-hop question and several edited facts, every of which can alter the ground-truth answer to the question.
4.1 Annotation Mistakes in One-Third Instances of MQuAKE
@@
Zhong et al. (2023) suggests using a randomly sampled 3,000 instance subset of MQuAKE to do the evaluation, which reduces the experimental costs, known as MQuAKE-3k. There are two issues of using MQuAKE-3k to evaluate KE. The first issue is that the new knowledge from different instances MQuAKE-3k can cause conflicts and mistakes in the ground-truth answers. In other words, the ground-truth answer from instance $A$ can be altered by the new knowledge from another instance $B$. We show an example of knowledge conflicts in MQuAKE-3k in Figure 5. These knowledge conflicts will make the ground-truth answers incorrect given the conflicted new knowledge because the inference on every instance could retrieve the new knowledge from all instances. We observe that 998 instances’ ground-truth answers are broken by the new knowledge conflicts across instances.
@@ 4.2 New Benchmark MQuAKE-2002 for More Precise Evaluation
To address the issue of annotation mistakes in MQuAKE-3k, we provide a new KE benchmark based on MQuAKE, which does not have any knowledge conflict across instances. This benchmark includes 2,002 instances, so we term it as MQuAKE-2002. We filter out the instances of which the ground-truth answers are broken by the new knowledge from other instances to produce MQuAKE-2002. Compared with MQuAKE-3k, our MQuAKE-2002 provides a more precise evaluation for KE methods, since it removes the annotation mistakes due to knowledge conflicts across instances. The data statistics of MQuAKE-2002 is provided in Table 2.
4.3 New Benchmark MQuAKE-hard for Challenging Evaluation
The second issue of MQuAKE-3k is that more than 60% instances in it only contain at most two edited facts that influence the answers, which are not challenging enough to evaluate the KE methods on handling multiple edited facts that can alter the ground-truth answers. We construct a more challenging subset of MQuAKE by selecting the instances that contain the highest number of edited facts per instance excluding MQuAKE-3k. We term this challenging set as MQuAKE-hard, which includes 429 instances and every instance contains four edited facts that can change the ground-truth answer. The data statistics of MQuAKE-hard is also provided in Table 2.
Benchmark # Instances # Hops per Instance # Edited Facts per Instance # Conflicted Instances MQuAKE-3k Zhong et al. (2023) 3,000 3.0 2.0 998 MQuAKE-2002 (Ours) 2,002 2.7 2.2 0 MQuAKE-hard (Ours) 429 4.0 4.0 0
@@ 5 Experiments
In this section, we evaluate the KE performance of our DeepEdit method when applied to LLMs, and compare against strong KE baselines Zhong et al. (2023). Our experimental settings closely follow those of the previous work Zhong et al. (2023) to ensure a fair comparison.
@@ 5.1 Experimental Settings
We take the following KE benchmarks MQuAKE-3k Zhong et al. (2023), MQuAKE-2002, and MQuAKE-hard for evaluation. We follow the existing work Zhong et al. (2023) to use accuracy as the main evaluation metric. For every instance, we test the accuracy of the first question in the instance, since different questions in the same instance are semantically equivalent and correspond to the same ground-truth answer. We follow Zhong et al. (2023) to set the KE batch size as 100 and full batch for evaluation. The KE batch size means the number of instances that provide new knowledge for retrieval when answering every question.
@@
Compared methods. We take the strong KE methods into comparison: MEND Mitchell et al. (2021), MEMIT Meng et al. (2022b), IKE Zheng et al. (2023), ICE Cohen et al. (2023), and MeLLo Zhong et al. (2023). MEND produces weight updates by transforming the raw fine-tuning gradients given an edited fact. MEMIT updates feedforward networks in a range of layers to encode the edited facts. IKE and ICE conduct KE by adding new knowledge to the input prompts. MeLLo is the state-of-the-art KE method that prompts the LLMs to generate subquestions and append new knowledge to the prompts. The former two KE methods are only applicable to white-box LLMs, while the latter three can fit the block-box LLMs.
@@
Method MQuAKE-3k MQuAKE-hard MQuAKE-2002 100 edited all edited 100 edited all edited 100 edited all edited LLaMA2-7B-Chat w/ MEND Mitchell et al. (2021) 7.2 3.9 0.5 0.2 7.7 4.1 LLaMA2-7B-Chat w/ MEMIT Meng et al. (2022b) 7.9 4.2 0.9 0.5 8.3 4.8 LLaMA2-7B-Chat w/ IKE Zheng et al. (2023) 12.5 6.2 1.2 0.5 12.9 6.5 LLaMA2-7B-Chat w/ ICE Cohen et al. (2023) 12.9 6.3 1.4 0.7 13.2 6.5 LLaMA2-7B-Chat w/ MeLLo Zhong et al. (2023) 16.3 10.8 1.9 1.6 17.1 11.8 LLaMA2-7B-Chat w/ DeepEdit (Ours) 27.6 11.2 14.5 7.0 29.5 12.9 GPT-3.5-Turbo-Instruct w/ IKE Zheng et al. (2023) 15.5 8.3 1.6 1.2 17.8 11.0 GPT-3.5-Turbo-Instruct w/ ICE Cohen et al. (2023) 16.1 8.5 2.1 1.4 18.1 11.5 GPT-3.5-Turbo-Instruct w/ MeLLo Zhong et al. (2023) 29.9 20.0 3.7 1.6 31.2 25.1 GPT-3.5-Turbo-Instruct w/ DeepEdit (Ours) 47.2 38.0 51.3 48.0 60.1 53.7 Text-Davinci-003 w/ IKE Zheng et al. (2023) 18.2 8.9 1.9 1.2 19.1 10.2 Text-Davinci-003 w/ ICE Cohen et al. (2023) 19.0 8.6 2.3 1.2 19.5 10.8 Text-Davinci-003 w/ MeLLo Zhong et al. (2023) 32.2 25.6 4.7 2.3 33.5 27.8 Text-Davinci-003 w/ DeepEdit (Ours) 49.1 38.8 52.0 49.2 60.9 54.8
@@
Table 3: Experimental results (accuracy; %) on the dataset MQuAKE-3k and MQuAKE-hard. “$k$ edited” denotes the KE batch size, which means the number of instances providing new knowledge on answering every question. Models with are closed source. The best KE result on every LLM is highlighted in bold font.
LLM as Constraint Verfier Conciseness Coherence Receptiveness Pertinence LLaMA2-7B-Chat 82.3 77.9 86.3 78.2 GPT-3.5-Turbo-Instruct 92.8 91.5 92.2 86.7
@@
Table 4: Experimental results (verification accuracy; %) of our verifiers for different constraints.
@@
Model configuration. We set the hyper-parameters of the baseline methods as suggested by their papers. We apply the KE methods to the popular LLMs GPT-3.5-Turbo-Instruct, Text-Davinci-003, and LLaMA2-7B-Chat OpenAI (2022); Touvron et al. (2023) for evaluation. The former two are black-box LLMs and the latter one is a white-box LLM. By default, we set the decoding temperature as 0.0 to minimize the randomness of LLMs’ outputs. Our constraint verifiers use the same LLM as the generator.
5.2 Overall Performance
We incorporate DeepEdit with several choices of LLMs to test its effectiveness on editing LLMs’ knowledge. We report the test accuracy on MQuAKE, MQuAKE-3k, MQuAKE-2002 in Table 3. When the KE batch size is 100, our DeepEdit improves the accuracy of question answering with KE by 69%, 58%, and 52% over MeLLo on the GPT-3.5-Turbo-Instruct, Text-Davinci-003, LLaMA2-7B-Chat respectively over the strong baseline method MeLLo on MQuAKE. On MQuAKE-hard, our DeepEdit outperforms the strong baseline method MeLLo by $\times$7, $\times$12, and $\times$10 times in terms of question answering accuracy with new knowledge on the LLMs LLaMA2-7B-Chat, Text-Davinci-003, and GPT-3.5-Turbo-Instruct respectively.
When the KE batch is full batch, our DeepEdit still outperforms the baseline methods and improves the KE accuracy by a significant margin. Our DeepEdit enhances the LLMs’ reasoning on new knowledge through higher conciseness, coherence, pertinence, and receptiveness. The above advantages are especially significant when the LLMs have to utilize multiple new knowledge to get the correct answers because more new knowledge requires LLMs to be correctly aware of the new knowledge on more reasoning steps. The correct knowledge incorporation with our decoding constraints significantly improves the KE performance given more new knowledge. Overall, our DeepEdit achieves substantial improvements on KE for LLMs and consistently outperforms the baseline KE methods on various benchmarks.
GPT-3.5-Turbo-Instruct w/ DeepEdit 49.1 $\Delta$ w/o Conciseness 46.1 $\downarrow 6\%$ w/o Coherence 42.9 $\downarrow 13\%$ w/o Receptiveness 45.2 $\downarrow 8\%$ w/o Pertinence 46.6 $\downarrow 5\%$
Table 5: Experimental results (accuracy; %) on the dataset MQuAKE of DeepEdit with KE batch size of 100.
5.3 Accuracy of Constraint Verification
We conduct an in-depth analysis on the constraint verification of our constraints. We randomly sample 4,000 pairs of sentences from MQuAKE that satisfy our constraints as positive samples and the same number of sentence pairs from different instances as the negative samples. We evaluate different verifiers based on LLaMA2-7B-Chat and GPT-3.5-Turbo-Instruct on the above samples. We report the verification accuracy in Table 4. We observe high verification accuracy of around 80% and over 90% on different constraints achieved by LLaMA2-7B-Chat and GPT-3.5-Turbo-Instruct respectively. This result matches the performance gap between two LLMs on general reasoning tasks OpenAI (2022); Touvron et al. (2023).
5.4 Efficiency of Depth-first Search
As analyzed in Section 3.3, we utilize the depth-first search to efficiently increase the LLMs’ reasoning depth with new knowledge. We evaluate the efficiency of our depth-first search using the 3-hop and 4-hop questions in the MQuAKE-3k benchmark, which holds the longest reasoning chains among the instances.
The methods we evaluate include: MeLLo, the breadth-first search-based decoding method introduced in Sec. 3.3, and our depth-first earch. The average number of reasoning steps and the time of text generation on every instance taken by different methods are reported in Figure 6. where the generation time is on a Linux Server using three A6000 GPUs.
We notice that, compared with the breadth-first search method, our DeepEdit significantly reduces the number of reasoning steps and the generation time. Our DeepEdit efficiently finds the reasoning chain with new knowledge by considering the importance of different step candidates. These results agree with the theoretical analysis in Sec. 3.3, since our DeepEdit’ time complexity is $\mathcal{O}(h)$ in the base case, while the time complexity of the breadth-first search is $\mathcal{O}(c^{h})$, much higher than our DeepEdit. Compared with the baseline MeLLo, our DeepEdit leads to much less generation time because our output text is purely the reasoning steps without the redundant subquestions and edited facts kept by MeLLo.
Figure 6: Inference costs of the baseline method MeLLo Zhong et al. (2023), breadth-first search introduced in Sec. 3.3, and our depth-first search.
5.5 Ablation Study
We conduct ablation studies on DeepEdit to empirically examine the contribution of its main technical components. We evaluate the influence of different constraints on DeepEdit. We report the experimental results of the ablation study of constraints in Table 5. We observe that removing each constraint leads to substantial performance degradation in different degrees. This indicates that different constraints contribute to more effective KE from the perspectives of progress, relevance, coherence, and awareness as a whole, and removing any of them will lead to more mistakes in the reasoning of LLMs. The above results also validate that the LLMs’ outputs cannot automatically satisfy the constraints that represent the desired properties without our DeepEdit. More experiments and case studies can be found in the appendix.
6 Conclusion
We explore a new paradigm for knowledge editing of black-box LLMs: decoding with constraints. We have designed a new decoding method DeepEdit to improve the knowledge editing of LLMs. DeepEdit enhances LLM’s outputs to soundly and efficiently incorporate new knowledge in the multi-step reasoning. In addition to DeepEdit, we provide two new benchmarks: MQuAKE-2002 and MQuAKE-hard to provide more precise and challenging evaluations for knowledge editing methods. Extensive experiments demonstrate the effectiveness of our DeepEdit. Future work includes exploring other constraints that can improve knowledge editing for large language models.
Acknowledgements
This work was partially supported by a DARPA ANSR program FA8750-23-2-0004. The views and conclusions are those of the authors and should not reflect the official policy or position of DARPA or the U.S. Government.
Appendix A Analysis on Knowledge Editing Batch Size
We follow the existing work Zhong et al. (2023) to evaluate the influence of the KE batch sizes on the performance. A larger batch size leads to the more new knowledge collected from more instances for retrieval. We present the test accuracy on MQuAKE-3k with different batch sizes of KE in Figure 7. We observe that the performance of GPT-3.5-Turbo-Instruct with MeLLo and our DeepEdit decreases with a larger batch size of knowledge editing. The reason is that with a larger batch, the difficulty of retrieving the relevant new knowledge to the target questions becomes higher. On the other hand, our DeepEdit still exhibits significant and consistent improvements over the strong KE method MeLLo.
Figure 7: Test accuracy (%) of question answering with new knowledge with different KE batch sizes. We present the performance of MeLLo and our DeepEdit applied on GPT-3.5-Turbo-Instruct.
@@ Appendix A Analysis on Knowledge Editing Batch Size
Appendix B Analysis on the Data where the Edited Facts do not Change the Ground-Truth Answers
Knowledge editing methods enhance the LLMs’ reasoning capability on the new knowledge. However, there exists the risk that the KE methods can degrade the LLMs’ performance on the data where no new knowledge changes their ground-truth answers. To evaluate whether our method DeepEdit degrades the performance of LLMs in this case. We do the following experiments, for every question-answering instance in MQuAKE-3k, we retrieve the edited facts only from other instances as the new knowledge, which would not change the ground-truth answer for the target question. In this setting, we first test the GPT-3.5-Turbo-Instruct without any KE methods, and compare the performance of GPT-3.5-Turbo-Instruct with the KE methods MeLLo and our DeepEdit. We present the performance in Table 6.
Method MQuAKE-3k GPT-3.5-Turbo-Instruct 55.7 GPT-3.5-Turbo-Instruct w/ MeLLo Zhong et al. (2023) 54.2 GPT-3.5-Turbo-Instruct w/ DeepEdit (Ours) 55.5
Table 6: Experimental results (accuracy; %) on the dataset MQuAKE-3k with no edited facts change the ground-truth answers.
The empirical results show that our DeepEdit exhibits higher robustness to give the correct answers for the multi-hop question answering than MeLLo. The performance drop of GPT-3.5-Turbo-Instruct with our DeepEdit when no edited facts change the ground-truth labels is negligible.
In addition to the above experiments, we also follow existing work Meng et al. (2022a); Zheng et al. (2023) to evaluate the Specificity scores of different knowledge editing methods on CounterFact given the KE batch size as full batch, which measures the accuracy of querying neighborhood prompts by Neighborhood Scores. In CounterFact, the neighborhood prompts share the same original object with the target prompt and these facts are not supposed to be edited. The results are presented in Table 7.
Method CounterFact LLaMA2-7B-Chat 85.3 LLaMA2-7B-Chat w/ MEND Mitchell et al. (2021) 57.8 LLaMA2-7B-Chat w/ MEMIT Meng et al. (2022b) 80.9 LLaMA2-7B-Chat w/ MeLLo Zhong et al. (2023) 82.7 LLaMA2-7B-Chat w/ DeepEdit (Ours) 84.1
Table 7: Specificity Score on the dataset CounterFact.
Our DeepEdit outperforms all the baseline KE methods in terms of the Specificity Score, which validates that our DeepEdit causes less performance degradation for LLMs than the baseline methods when the edited facts do not change the ground-truth labels.
Appendix C Iteration-wise Visualization of Our DeepEdit
We present the iteration-wise outputs of our DeepEdit on an instance in Figure 8.
Appendix D Case Study